OpenStudy (anonymous):

If iz^3 + z^2 -z + i =0 then |z| is?

OpenStudy (anonymous):

ok lets see what we have here

OpenStudy (anonymous):

man this is not that easy ... lol :)

OpenStudy (anonymous):

\[iz^3 + z^2 -z + i =0\]\[iz^3 -i^2 z^2 -z + i =0\]\[iz^2(z-i) -(z - i) =0\]\[(z - i)(iz^2-1) =0\]

OpenStudy (anonymous):

For all the roots |z|=1

OpenStudy (unklerhaukus):

mukushla, how did you know how to do that ?

OpenStudy (anonymous):

and...u r correct @UnkleRhaukus can u explain this

OpenStudy (unklerhaukus):

the factored form shows \[(z-i)=0\] \(z=i\) is a solution _____ \[|z|=|i|\]

OpenStudy (unklerhaukus):

the other solution\[(iz^2−1)=0\] \[iz^2=1\] \[z=\sqrt{\frac1{i}}=\sqrt{-i}=i\sqrt i\]

OpenStudy (anonymous):

Same result if you take the real part, then the imaginary part of both sides of the original equation.

OpenStudy (anonymous):

\[iz^2−1=0\]\[z^2=-i\]\[|z^2|=|z|^2=1\]\[|z|=1\]

OpenStudy (anonymous):

lol...@mukushla there will be two soln for ur eq

OpenStudy (anonymous):

for |z| ? what are the solutions?

OpenStudy (unklerhaukus):